# Dynamical Logical Qubits in the Bacon-Shor Code

### v1

@@ Abstract
The Bacon-Shor code is a quantum error correcting subsystem code composed of weight 2 check operators that admits a single logical qubit, and has distance $d$ on a $d\times d$ square lattice. We show that when viewed as a Floquet code, by choosing an appropriate measurement schedule of the check operators, it can additionally host several dynamical logical qubits. Specifically, we identify a period 4 measurement schedule of the check operators that preserves logical information between the instantaneous stabilizer groups. Such a schedule measures not only the usual stabilizers of the Bacon-Shor code, but also additional stabilizers that protect the dynamical logical qubits against errors. We show that the code distance of these Floquet-Bacon-Shor codes scales as $\Theta(d/\sqrt{k})$ on a $d\times d$ lattice with $k$ dynamical logical qubits, along with the logical qubit of the parent subsystem code. Moreover, several errors are shown to be self-corrected purely by the measurement schedule itself.
@@ I Introduction
The original honeycomb Floquet code [4] was defined on a hexagonal torus, while generalizations to 3-colorable and 3-valent planar graphs were explored in [6]. Due to a non-trivial automorphism between electric and magnetic operators across measurement cycles, adding boundaries to the model is non-trivial, but were introduced in [7]. Benchmarking studies [8, 9, 10] showed a competitive threshold for the model, with at least one small scale experiment[11] demonstrating stabilizer measurements. Since then, several examples of Floquet codes have been constructed, including those without parent subsystem codes[12], those involving the color code[13, 14], constructions based on hyperbolic geometries[15, 16], in 3D Euclidean space based on fracton order[17], rewinding the measurement schedule[18], as well as introducing twists in the geometry[19]. Generalizations of Floquet code constructions have also been explored using the ZX calculus[20, 21], adiabatic paths[22], anyon condensation[13], as well as aperiodic measurement schedules in the form of dynamic automorphism codes[23].
@@
While the dynamics of previously constructed Floquet codes typically requires a description of the embedded toric codes in its ISGs, and may perhaps be intuitively understood through adiabatic paths [22] or anyon condensation [13], the construction of the Floquet code we describe here follows directly from the subsystem structure of the parent Bacon-Shor code, without the need to move to a condensed matter perspective. In this sense, it hopefully offers a comparatively simpler example of a Floquet code. It also comes naturally equipped with a boundary, without the complication of non-trivial automorphisms between electric and magnetic operators [4, 7, 6]. Furthermore, this Floquet code is naturally defined on a square lattice, whose simple lattice structure may be appealing from an experimental perspective. To our knowledge, the model introduced here provides a first example of a Floquet code on a square lattice, though it has been noted that a square lattice construction may also be possible using an alternative approach[24]. It is an example of a CSS Floquet code[12, 13], since all the stabilizers are of either $X$ or $Z$ type. Moreover, unlike the original honeycomb code[4], it also serves as an example of a Floquet code whose parent subsystem code hosts a non-zero number of logical qubits, as well as one which may host several dynamical logical qubits. Perhaps most importantly, it demonstrates the construction of a Floquet code using essentially only the subsystem structure of the parent code through the addition of gauge defects. We strongly suspect that this approach may be much more widely applicable in constructing more examples of Floquet codes.
@@ III Floquet-Bacon-Shor Code
In order to measure the stabilizers of the Bacon-Shor subsystem code, it suffices to use a period 2 measurement schedule, such that we measure all the horizontal $XX$ checks in the first round, followed by all the vertical $ZZ$ checks. From a Floquet perspective, this corresponds to periodically moving between two different gauge fixings of the Bacon-Shor code, one where the $\overline{X}$ operators of all the gauge qubits has been fixed, followed by fixing all the $\overline{Z}$ gauge qubit operators. We can view either gauge fixing as a stabilizer code that contains the Bacon-Shor stabilizers as a subset, but also includes the $\overline{X}$ or $\overline{Z}$ gauge qubit operators. It does not matter whether these gauge qubit operators are fixed to ±$\pm 1$, and in general they will take random values upon each measurement, with the only constraint being the values of the stabilizers.
In order to free up space for an additional logical qubit, we must unfix one of the gauge degrees of freedom. In general, such an unfixing, or gauge defect, is difficult to maintain while ensuring all the Bacon-Shor stabilizers also get measured. However, as we show here, it is possible to maintain such a gauge defect in a dynamical sense, such that the encoded space for the corresponding logical qubit changes across measurement rounds, similar to the examples of previously constructed Floquet codes.
@@ III.1 Measurement schedule and ISGs
In particular, we can adopt the period 4 measurement schedule shown in the top part of Fig. 2. The core idea here is that we distribute the measurement of the Bacon-Shor stabilizers into four, instead of two, rounds, allowing us to prevent one of the gauge degrees of freedom from being fixed, thereby introducing a gauge defect that evolves across the measurement rounds and serves as the dynamical logical degree of freedom.
In order to describe the measurement process, it is helpful to label four of the square plaquettes that serve a special role in identifying the dynamical logical qubit. For simplicity, we will assume a $d\times d$ square lattice. As in Section II, we denote the location of square plaquettes by the location of their top-right corner. Since square plaquettes are in one-to-one correspondence with gauge qubits, we will use the two terms interchangeably. For odd $d$, the gauge qubit A is identified as the plaquette at location $\left(\frac{d-1}{2},\frac{d+1}{2}\right)$, B at $\left(\frac{d+1}{2},\frac{d+1}{2}\right)$, C at $\left(\frac{d+1}{2},\frac{d-1}{2}\right)$, and D at $\left(\frac{d-1}{2},\frac{d-1}{2}\right)$. Similarly, for even $d$, A is identified as the plaquette at $\left(\frac{d}{2}-1,\frac{d}{2}+1\right)$, B at $\left(\frac{d}{2},\frac{d}{2}+1\right)$, C at $\left(\frac{d}{2},\frac{d}{2}\right)$, and D at $\left(\frac{d}{2}-1,\frac{d}{2}\right)$. In either case, these four plaquettes all share a single corner that lies roughly at the center of the square lattice. We also denote columns AD and BC, and rows AB and CD as the unique columns and rows containing the obvious choice of plaquettes, as well as edges AB, BC, CD and AD as the edges between the respective plaquettes.
@@
In order to preserve logical information across these four ISGs, it is also necessary that some representative, upto instantaneous stabilizer operations, of a logical operator at round $r$ persists as a logical operator in the next round $r+1$. This is formalized in Eq. (2). Roughly speaking, this condition ensures that if |$\ket{\psi^{(r)}}=\alpha\ket{0^{(r)}}+\beta\ket{1^{(r)}}$ is the state of the dynamical logical qubit at round $r$, then the measurement schedule projects it to |$\ket{\psi^{(r)}}\rightarrow\alpha\ket{0^{(r+1)}}+\beta\ket{1^{(r+1)}}=\ket{\psi^{(r+1)}}$ at round $r+1$, where the codewords |$\ket{0^{(r)}}$ etc change across the ISGs, but the pair of coefficients $\alpha$ and $\beta$, and therefore the logical information, does not. Specifically, this condition says that if $S^{A}$ and $S^{B}$ are two stabilizer groups associated with stabilizer codes $A$ and $B$ respectively, with associated logical operators $O^{A}$ and $O^{B}$, where $O^{j}\subset\mathcal{N}(S^{j})\setminus S^{j}$, then if there exist $s^{A}\in S^{A}$ and $s^{B}\in S^{B}$ such that
@@ III.2 Multiple Dynamical Qubits
Introducing additional dynamical logical degrees of freedom amounts to adding more gauge defects to the measurement schedule, and is straightforward to generalize from the construction above for a single dynamical logical qubit. Consider relabeling the set of plaquettes $(A,B,C,D)$ described above as $(A_{0},B_{0},C_{0},D_{0})$, and identifying another such similar set of plaquettes $(A_{1},B_{1},C_{1},D_{1})$ where we repeat the measurement schedule identified above. Thus, for instance, we would now have two defect columns, $A_{0}D_{0}$ and $A_{1}D_{1}$, at round 0. The code distance would now be proportional to the shorter of the two distances from any edge of the square lattice to either of the two resultant gauge defects. To make one of these distances the same for both gauge defects, we can put them in the same plaquette column or row. Suppose, without loss of generality, that we put them in the same row. The optimal choice would be to place both gauge defects roughly $d/3$ distance away from each other, as well as either vertical edge of a $d\times d$ square lattice, but roughly $d/2$ away from the horizontal edges. Since this already reduces the code distance to $\sim d/3$, one might as well add two more gauge defects such that all of these lie roughly $d/3$ units away from any edge of a $d\times d$ square lattice.
@@ IV Error Detection and Correction
Here, we assume an error model of independent $X$ and $Z$ errors on all the qubits between measurement rounds, and independent measurement errors of the $XX$ and $ZZ$ check operators. A similar error model was assumed in [4], but there the model also assumed perfect measurements of the check operators, with measurement errors of e.g. $XX$ checks being equivalent to a single $Z$ error on one of the qubits supported by this check before and after the measurement of this check operator. In our case, such an error pattern would flip two checks supported by that qubit, and would make it inequivalent to a measurement error on a single check. Since $X$ operators can be commuted past the measurements of rounds 0 and 2, we assume that $X$ errors only occur between either rounds 0 and 1, or between rounds 2 and 3. Similarly, $Z$ errors can be commuted past the measurements of round 1 and 3, and we assume that these occur either between rounds 1 and 2, or between rounds 3 and 0 (mod 4).
The decoding graph is typically defined as the graph whose vertices are given by the stabilizer measurements, and edges are given by independent errors, or faults. In the Floquet-Bacon-Shor construction, described in previous sections of the paper, we can have independent errors that flip up to 4 stabilizers. Thus, we are naturally led to the slight generalization of a decoding hyper-graph. A hyper-graph is a generalization of a graph where instead of an edge connecting two vertices, we have a hyper-edge that connects $n$ vertices. In the decoding hyper-graph, a hyper-edge connecting 4 vertices would correspond to precisely such an independent error. This makes decoding more complicated in principle, as standard decoding algorithms involving Minimum Weight Perfect Matching (MWPM) or Union Finding (UF) typically work with edges, not hyper-edges. However, some recent work [29] has shown that one could nevertheless adapt such techniques to work with decoding hyper-graphs as well.
@@
In the Floquet-Bacon-Shor code described above, we have additional stabilizers that protect the encoded dynamical logical qubits. These are given by the gauge fixed operators that lie along the defect columns and rows of the various ISGs. We will refer to these as transient stabilizers, to distinguish them from the usual Bacon-Shor stabilizers that run along the length of the entire lattice, which we shall refer to as permanent stabilizers. The collection of such gauge operators, or transient stabilizers, in any defect column or row form a domain wall across which single qubit errors cannot propagate, since we skip the measurements along that column or row. For instance, assuming temporarily a single dynamical logical qubit for simplicity, at round $4k$ ($k\geq 1$), we do not measure any of the $XX$ checks in the defect column AD except on the edge AD, so that single qubit $X$ errors are equivalent, up to instantaneous stabilizer operations, to any single $X$ error on either side of the domain wall in the column AD, depending on which side of this wall the error occurred, except along the row that contains the edge AD. This means that unlike the Bacon-Shor code, an even number of $X$ errors, for instance, can bring us out of the dynamical codespace if we have an odd number of $X$ errors on both sides of the domain wall.
Measurement errors of the permanent stabilizers are given in the decoding hyper-graph as edges connecting the same stabilizer across consecutive time slices. A defect is referred to as a flip in the recorded value of a stabilizer. A measurement error of a permanent stabilizer will cause two defects in the decoding hyper-graph, once due to the erroneous flip in recorded value, and the second when its value flips back to the correct one, and can therefore be decoded by standard matching algorithms. In the case of measurement errors of transient stabilizers, one has to be a bit more careful. Suppose we have just carried out the measurements of round 1 mod 4, and one of the $Z$-type check operators along some plaquette row sufferred a measurement error. If it occurred to the left of column AD, then it flips only the value of the permanent $Z$-type stabilizer along this row. If it occurs on the check between cells A and B, then it flips the value of permanent $Z$-type stabilizer along this row, as well as the transient stabilizer given by the gauge fixed $\overline{Z}$ operator of the gauge qubit lying at the intersection of this row and column AD. If it instead occurs on any check further right along this row, then it flips the previous two stabilizers, as well as the gauge fixed $\overline{Z}$ operator of the gauge qubit at the intersection of this row and column BC. Thus, measurement errors of transient stabilizer in general will lead to hyper-edges connecting several vertices in the decoding hyper-graph.

### v2

@@ Abstract
The Bacon-Shor code is a quantum error correcting subsystem code composed of weight 2 check operators that admits a single logical qubit, and has distance $d$ on a $d\times d$ square lattice. We show that when viewed as a Floquet code, by choosing an appropriate measurement schedule of the check operators, it can additionally host several dynamical logical qubits. Specifically, we identify a period 4 measurement schedule of the check operators that preserves logical information between the instantaneous stabilizer groups. Such a schedule measures not only the usual stabilizers of the Bacon-Shor code, but also additional stabilizers that protect the dynamical logical qubits against errors. We show that the code distance of these Floquet-Bacon-Shor codes scales as $\Theta(d/\sqrt{k})$ on an $n=d\times d$ lattice with $k$ dynamical logical qubits, along with the logical qubit of the parent subsystem code. Unlike the usual Bacon-Shor code, the Floquet-Bacon-Shor code family introduced here can therefore saturate the subsystem bound $kd=O(n)$. Moreover, several errors are shown to be self-corrected purely by the measurement schedule itself. This work provides insights into the design space for dynamical codes and expands the known approaches for constructing Floquet codes.
@@ I Introduction
The original honeycomb Floquet code [4] was defined on a hexagonal torus, while generalizations to 3-colorable and 3-valent planar graphs were explored in [6]. Due to a non-trivial automorphism between electric and magnetic operators across measurement cycles, adding boundaries to the model is non-trivial, but were introduced in [7]. Benchmarking studies [8, 9, 10, 11] showed a competitive threshold for the model, with at least one small scale experiment[12] demonstrating stabilizer measurements. Since then, several examples of Floquet codes have been constructed, including those without parent subsystem codes[13], those involving the color code[14, 15], constructions based on hyperbolic geometries[16, 17], in 3D Euclidean space based on fracton order[18], involving qudits [19], using a path integral approach [20, 21], rewinding the measurement schedule[22], using twists in the geometry[23], as well as those tailored for biased noise [24]. Generalizations of Floquet code constructions have also been explored using the ZX calculus[25, 26, 27, 28], adiabatic paths[29], anyon condensation[14], as well as aperiodic measurement schedules in the form of dynamic automorphism codes[30].
@@
A main aim of this work is to provide insights into the design space for dynamical codes. While the dynamics of previously constructed Floquet codes typically requires a description of the embedded toric codes in its ISGs, and may perhaps be intuitively understood through adiabatic paths [29] or anyon condensation [14], the construction of the Floquet code we describe here follows directly from the subsystem structure of the parent Bacon-Shor code, without the need to move to a condensed matter perspective. In this sense, it provides a complementary approach to the design of dynamical codes and offers a comparatively simpler example of a Floquet code. This code also comes naturally equipped with a boundary, without the complication of non-trivial automorphisms between electric and magnetic operators [4, 7, 6]. Furthermore, this Floquet code is naturally defined on a square lattice which may be appealing from an experimental perspective, and has a period 4 measurement schedule. A Floquet code on a square lattice was previously constructed in [15] by transforming the 2D hexagonal color code [31] to a square lattice geometry using the ZX calculus [32], resulting in a period 13 measurement schedule involving 1 and 2 body checks, mapping detectors and logical operators of the original color code to those of the new code. Furthermore, square lattice construction may be possible using an alternative approach [33]. In the code described here, the square lattice construction appears more simply and naturally from the subsystem structure of the parent Bacon-Shor code, with the stabilizers and logical operators being readily identifiable. It is an example of a CSS Floquet code[13, 14], since all the stabilizers are of either $X$ or $Z$ type. Moreover, unlike the original honeycomb code[4], it also serves as an example of a Floquet code whose parent subsystem code hosts a non-zero number of logical qubits, which can then be augmented by our construction to host several dynamical logical qubits. Perhaps most importantly, it demonstrates the construction of a Floquet code using essentially only the subsystem structure of the parent code through the addition of gauge defects. We strongly suspect that this approach may be more widely applicable in constructing more examples of Floquet codes.
@@ III Floquet-Bacon-Shor Code
In order to measure the stabilizers of the Bacon-Shor subsystem code, it suffices to use a period 2 measurement schedule, such that we measure all the horizontal $XX$ checks in the first round, followed by all the vertical $ZZ$ checks. From a Floquet perspective, this corresponds to periodically moving between two different gauge fixings of the Bacon-Shor code, one where the $\overline{X}$ operators of all the gauge qubits has been fixed, followed by fixing all the $\overline{Z}$ gauge qubit operators, as depicted in Fig. 2. We can view either gauge fixing as a stabilizer code that contains the Bacon-Shor stabilizers as a subset, but also includes either the $\overline{X}$ or $\overline{Z}$ gauge qubit operator. It does not matter whether these gauge qubit operators are fixed to ±$\pm 1$, and in general they will take random values upon each measurement, with the only constraint being the values of the stabilizers.
Figure 2: A simple period 2 measurement schedule (top) for the Bacon-Shor subsystem code, given by measuring all the horizontal $XX$ checks (colored red) and therefore all the $X$-type (vertical) stabilizers, followed by measuring all the vertical $ZZ$ checks (colored blue) and similarly all the $Z$-type (horizontal) stabilizers. From a Floquet perspective (bottom), this corresponds to moving between two instantaneous stabilizer groups (ISGs) that are total gauge fixings of the parent subsystem code, where we fix either all the $X$ operators of the gauge qubits in the first round, or all their $Z$ operators in the second round.
In order to free up space for an additional logical qubit, we must unfix one of the gauge degrees of freedom. In general, such an unfixing, or gauge defect, is difficult to do while ensuring all the Bacon-Shor stabilizers also get measured. However, as we show here, it is possible to maintain such a gauge defect in a dynamical sense, where the encoded space for the corresponding logical qubit changes across measurement rounds, as seen in previously constructed Floquet codes.
@@ III.1 Measurement schedule and ISGs
We can adopt the period 4 measurement schedule shown in the top part of Fig. 3. The core idea here is that we distribute the measurement of the Bacon-Shor stabilizers into four, instead of two, rounds, allowing us to prevent one of the gauge degrees of freedom from being fixed, thereby introducing a gauge defect that evolves across the measurement rounds and serves as the dynamical logical degree of freedom.
In order to describe the measurement process, it is helpful to label four of the square plaquettes that serve a special role in identifying the dynamical logical qubit. For simplicity, we will consider a $d\times d$ square lattice. As in Section II, we label the location of square plaquettes with the location of their top-right corner. Since square plaquettes are in one-to-one correspondence with gauge qubits, we will use the two terms interchangeably. For odd $d$, the gauge qubit A is identified as the plaquette at location $\left(\frac{d-1}{2},\frac{d+1}{2}\right)$, B at $\left(\frac{d+1}{2},\frac{d+1}{2}\right)$, C at $\left(\frac{d+1}{2},\frac{d-1}{2}\right)$, and D at $\left(\frac{d-1}{2},\frac{d-1}{2}\right)$. Similarly, for even $d$, A is identified as the plaquette at $\left(\frac{d}{2}-1,\frac{d}{2}+1\right)$, B at $\left(\frac{d}{2},\frac{d}{2}+1\right)$, C at $\left(\frac{d}{2},\frac{d}{2}\right)$, and D at $\left(\frac{d}{2}-1,\frac{d}{2}\right)$. In either case, these four plaquettes all share a single corner that lies roughly at the center of the square lattice. We also denote columns AD and BC, and rows AB and CD as the unique columns and rows containing the obvious choice of plaquettes, as well as edges AB, BC, CD and AD as the edges between the respective plaquettes.
@@
In order to preserve logical information across these four ISGs, it is necessary that some representative, up to instantaneous stabilizer operations, of a logical operator at round $r$ persists as a logical operator in the next round $r+1$. This is formalized in Eq. (2). Roughly speaking, this condition ensures that if |$\ket{\psi^{(r)}}=\alpha\ket{0^{(r)}}+\beta\ket{1^{(r)}}$ is the state of the dynamical logical qubit at round $r$, then the measurement schedule projects it to |$\ket{\psi^{(r)}}\rightarrow\alpha\ket{0^{(r+1)}}+\beta\ket{1^{(r+1)}}=\ket{\psi^{(r+1)}}$ at round $r+1$, where the codewords |$\ket{0^{(r)}}$ etc change across the ISGs, but the pair of coefficients $\alpha$ and $\beta$, and therefore the logical information, does not. Specifically, this condition says that if $S^{A}$ and $S^{B}$ are two stabilizer groups associated with stabilizer codes $A$ and $B$ respectively, with associated logical operators $O^{A}$ and $O^{B}$, where $O^{j}\subset\mathcal{N}(S^{j})\setminus S^{j}$, then if there exist $s^{A}\in S^{A}$ and $s^{B}\in S^{B}$ such that
@@ III.2 Multiple Dynamical Qubits
Introducing additional dynamical logical degrees of freedom amounts to adding more gauge defects to the measurement schedule. It is straightforward to generalize from the construction above for a single dynamical logical qubit. Consider relabeling the set of plaquettes $(A,B,C,D)$ described above as $(A_{0},B_{0},C_{0},D_{0})$, and identifying another such similar set of plaquettes $(A_{1},B_{1},C_{1},D_{1})$ where we repeat the measurement schedule identified above. Thus, for instance, we would now have two defect columns, $A_{0}D_{0}$ and $A_{1}D_{1}$, at round 0. The code distance would now be proportional to the shorter of the two distances from any edge of the square lattice to either of the two resultant gauge defects. To make one of these distances the same for both gauge defects, we can put them in the same plaquette column or row. Suppose, without loss of generality, that we put them in the same row. The optimal choice would be to place both gauge defects roughly $d/3$ distance away from each other, as well as either vertical edge of a $d\times d$ square lattice, but roughly $d/2$ away from the horizontal edges. Since this already reduces the code distance to $\sim d/3$, one might as well add two more gauge defects such that all of these lie roughly $d/3$ units away from any edge of a $d\times d$ square lattice.
@@
Note that the usual Bacon-Shor code defined on an $\sqrt{n}\times\sqrt{n}$ lattice with $n$ physical qubits has code distance $\sqrt{n}$ and admits a single logical qubit. Meanwhile, the bound for local 2D subsystem codes relating these parameters is $kd=O(n)$ [38]. In the case of the Floquet-Bacon-Shor code family introduced here, we can choose the number of dynamical gauge defects to be a constant fraction of the total number of physical qubits, which would also simultaneously set the code distance to be a constant. In other words, we would have $k=\Theta(n)$ and $d=\Theta(1)$, which does saturate the bound $kd=O(n)$, unlike the parent Bacon-Shor subsystem code.
@@ IV Error Detection and Correction
Here, we assume an error model of independent $X$ and $Z$ errors on all the qubits between measurement rounds, and independent measurement errors of the $XX$ and $ZZ$ check operators. A similar error model was assumed in [4], but there the model also assumed perfect measurements of the check operators, with measurement errors of e.g. $XX$ checks being equivalent to a single $Z$ error on one of the qubits supported by this check before and after the measurement of this check operator. In our case, such an error pattern would flip two checks supported by that qubit, and would make it inequivalent to a measurement error on a single check. Since $X$ operators can be commuted past the measurements of rounds 0 and 2, we assume that $X$ errors only occur between either rounds 0 and 1, or between rounds 2 and 3. Similarly, $Z$ errors can be commuted past the measurements of round 1 and 3, and we assume that these occur either between rounds 1 and 2, or between rounds 3 and 0 (mod 4). To ensure fault tolerance in the presence of measurement errors, we would need to repeat each measurement round $O(d)$ many times on a $d\times d$ lattice, not unlike the case of the usual Bacon-Shor code, but with the important difference that each round itself must be repeated $O(d)$ many times, and not the measurement cycle in which each measurement round occurs only once. This is so because we form detectors out of temporary stabilizers that take randomized measured values every measurement cycle, as we shall elaborate below. In the usual Bacon-Shor code, it would make no difference whether we repeated the rounds or the cycles, as all the stabilizers that are used to form detectors persist across all rounds.
@@
In the Floquet-Bacon-Shor code described above, we have additional stabilizers that protect the encoded dynamical logical qubits. These are given by the gauge fixed operators that lie along the defect columns and rows of the various ISGs. We will refer to these as transient stabilizers, to distinguish them from the usual Bacon-Shor stabilizers that run along the length of the entire lattice, which we shall refer to as permanent stabilizers. As we discuss below, a decoding graph can be built up entirely from transient stabilizers, so that standard matching decoding algorithms can be applied. The collection of such gauge operators, or transient stabilizers, in any defect column or row form a domain wall across which single qubit errors cannot propagate, since we skip the measurements along that column or row. For instance, assuming temporarily a single dynamical logical qubit for simplicity, at round $4k$ ($k\geq 1$), we do not measure any of the $XX$ checks in the defect column AD except on the edge AD, so that single qubit $X$ errors are equivalent, up to instantaneous stabilizer operations, to any single $X$ error on either side of the domain wall in the column AD, depending on which side of this wall the error occurred, except along the row that contains the edge AD. This means that unlike the Bacon-Shor code, an even number of $X$ errors, for instance, can bring us out of the dynamical codespace if we have an odd number of $X$ errors on both sides of the domain wall.